Soybean meal is 12% protein, cornmeal is 6% protien. How many pounds of each should be mixed together in order to get a 240 lb mixture that is 10% protien?
Answer by robertb(5830)

0.12s + 0.06(240 - s) = 0.10*240,
0.12s + 14.4 - 0.06s = 24,
0.06s = 9.6,
s = 160, the # lbs of soybean meal, and
240 - 160 = 80, the # lbs of cornmeal.